OpenStudy (anonymous):

How do i solve log(x^2+7)=3 plz show ur work

OpenStudy (reemii):

apply the exponential function on both sides, you obtain \(x^2+7=e^3\). you can solve this formula easily. (two solutions)

OpenStudy (amistre64):

umm, in lower math classes, log is base 10, not natural

OpenStudy (amistre64):

so this is either =10^3, or =e^3 :/

OpenStudy (reemii):

oh ok. then what you do is \[10^{\log_{10}(x^2+7)} = 10^3\\ x^2+7 = 10^3\]. thanks for the reminder

OpenStudy (amistre64):

real analysis gets into the "log" function as the natural log :)

OpenStudy (anonymous):

i am lost

OpenStudy (reemii):

we are wondering if your log is actually \(\log_{10}\) or \(\log_e\) . but the method is the same.

OpenStudy (anonymous):

i dont know my teacher gave me this got extrca cridet

OpenStudy (reemii):

if you don't kow, just say that you choose \(\log_{10}\), and do as in my msg starting with "oh ok". gl.

OpenStudy (anonymous):

ok tysvm